# here title TetraLoss: Improving the Robustness of Face Recognition against Morphing Attacks

### v1

@@ VII CONCLUSION
In this paper, we proposed a novel framework and loss function to make existing deep learning-based face recognition systems more robust against morphing attacks. To this end, a neural network architecture was optimised by a novel TetraLoss function to learn to separate morphing attack embeddings from embeddings of the subjects contributing to the morphed samples. The results show that the proposed method can seamlessly be applied on top of existing state-of-the-art face recognition systems to enhance their performance to morphing attacks while maintaining good face recognition performance. Specifically, using two distinct backbone architectures, the RIAPAR improves with at least 45% points for the operational relevant threshold at FMR=0.1% for ArcFace, MagFace, and AdaFace. A direct result of this work is that the TetraLoss can be employed to make biometric recognition systems robust against other digital or physical manipulations. As future work, we plan to dive into the generation of synthetic data to improve the generalisation capability of the proposed system across other attacks and environmental conditions.

### v2

@@ VII CONCLUSION
In this paper, we proposed a novel framework and loss function to make existing deep learning-based face recognition systems more robust against morphing attacks. To this end, a neural network architecture was optimised by a novel TetraLoss function to learn to separate morphing attack embeddings from bona fide embeddings extracted from face images of the subjects contributing to the morphed samples. The results show that the proposed method can seamlessly be applied on top of existing state-of-the-art face recognition systems to increase their resilience to morphing attacks while still achieving good face recognition performance on bona fide data. Specifically, using two distinct backbone architectures, the RIAPAR improves with at least 45% points for the operational relevant threshold at FMR=0.1% for ArcFace, MagFace, and AdaFace.